# A Survey on 3D Gaussian Splatting Applications: Segmentation, Editing, and Generation

### v2

@@ 4 Performance Comparison
TABLE III: Quantitative 3D interactive segmentation on NVOS [287] and SPIn-NeRF[67] in terms of mIoU and mAcc.
@@
TABLE IV: Performance comparison of open-vocabulary 3D semantic segmentation on ScanNet [280] benchmark. The evaluation metrics are mIoU and mAcc, respectively.
@@
4.1 Performance Benchmarking: 3DGS Segmentation
$\bullet$ Datasets. Table I summarizes the key features of the datasets for 3DGS segmentation, followed by detailed descriptions.
@@ 4.1 Performance Benchmarking: 3DGS Segmentation
$\bullet$ Results. Following the taxonomy in Sec.2.1.1, we report results across four representative segmentation settings. For 3D instance segmentation (Table II), Unified-Lift [20] achieves the best overall performance on both the Replica [285] and LERF-Mask [22] datasets, demonstrating the effectiveness of consistency-aware end-to-end learning. In the setting of 3D interactive segmentation (Table III), GaussianCut [29] performs best on NVOS [279], while Click-Gaussian [24] attains the highest mIoU on SPIn-NeRF [67]. Notably, GaussianCut is a training-free method that flexibly supports both 2D clicks and text prompts as input. For open-vocabulary 3D semantic segmentation (Table IV), PanoGS [112] and InstanceGaussian [31] achieve the best performance under point-based evaluation protocols, whereas CAGS [103] performs best under 3D Gaussian-based evaluation. In the open-vocabulary 2D semantic segmentation setting, VLGaussian [28] reaches a state-of-the-art mIoU of 97.1% on the 3D-OVS benchmark [63], significantly outperforming existing approaches in Table V. Table VI further presents results for two subtasks: open-vocabulary 2D localization (detection) and semantic segmentation. FMLGS [98] achieves the best performance on the segmentation task, while VLGaussian [28] leads in localization. Across these settings, most methods leverage powerful vision-language foundation models such as CLIP and SAM to enhance open-vocabulary capabilities.
TABLE V: Quantitative open-vocabulary 2D semantic segmentation on 3D-OVS [63] dataset in terms of mIoU.
@@
TABLE VI: Quantitative results on LERF-OVS [26] Dataset. We report the open-vocabulary localization accuracy (mAcc) and 2D semantic segmentation (mIoU).
@@
TABLE VII: Statistics of representative 3DGS editing datasets. See §4.2 for more detailed descriptions.
  Dataset Venue #Scenes #Views(Avg.) Characterization
TABLE VIII: Performance comparison of 3DGS editing on Mip-NeRF360 [281] and IN2N [75] datasets. The similarity for CLIP Text-Image Direction,CLIP Text-Image, CLIP Image-Image, DINO are denoted as $\text{CLIP}_{dir}$, $\text{CLIP}_{T2I}$, $\text{CLIP}_{I2I}$, and $\text{DINO}_{sim}$, respectively.
@@
TABLE IX: Performance comparison of 3DGS style transfer on Mip-NeRF360 [281] dataset, evaluated by LPIPS and RMSE.
@@ 4.2 Performance Benchmarking: 3DGS Editing
$\bullet$ Datasets. We summarize the key features of the datasets in Table VII, with detailed descriptions provided below.
@@
$\bullet$ Results. For 3D image editing tasks, the most commonly used datasets are Mip-NeRF360 [281] and IN2N [75]. We adopt these two datasets as benchmarks to record the performance of various methods. However, as most existing 3D editing methods are evaluated on different scene subsets, it remains difficult to perform a comprehensive and fair comparison under a unified protocol. Given that IN2N [75] is widely adopted as a baseline, we report the relative gains of each method over IN2N to enable an indirect yet relatively fair comparison, as summarized in Table VIII. Overall, GaussianVTON[155] demonstrates the best performance across multiple metrics. For the style transfer task, we use Mip-NeRF 360 [281] as the benchmark and compare the results of different methods in Table IX, where SemanticSplatSty.[167] achieves the best overall performance.
TABLE X: Statistics of representative 3DGS generation datasets. See §4.3 for more detailed descriptions.
  Dataset Venue #Type #Scenes Characterization
[…]
TABLE XI: Quantitative 3DGS generation on GSO [283] dataset. CLIP Image-Image Similarity as $\text{CLIP}_{I2I}$.
@@
TABLE XII: Quantitative 3DGS generation on Objaverse [282] dataset. CLIP Text-Image Similarity as $\text{CLIP}_{T2I}$.
@@ 4.3 Performance Benchmarking: 3DGS Generation
$\bullet$ Datasets. We summarize the key features of the datasets in Table X, with detailed descriptions provided below.
@@
Objaverse [282] is a large-scale dataset of objects with 800K+ (and growing) 3D models with descriptive captions, tags, and animations.
@@
$\bullet$ Results. We select two representative datasets, GSO [283] and Objaverse [282], as benchmarks for 3DGS generation. The performance of various methods is summarized in Table XI and Table XII, respectively. Due to the limited number of publicly comparable baselines, we provide a quantitative assessment based on the available results. On GSO dataset, NovelGS [217] achieves the best performance by tailoring the diffusion process to suit the 3D Gaussian representation. It supports both text-to-3D and image-to-3D generation, demonstrating strong generalization across input modalities. On Objaverse dataset, Atlas-Gaussians [207] performs best, leveraging latent space optimization to enhance generation quality. It primarily supports text-conditioned generation and shows promising results on open-domain content.
@@ 5 Future Directions
$\bullet$ Faithful Evaluation Metrics. Current evaluation protocols often inherit metrics from traditional 2D or NeRF-based tasks, which may not align well with 3DGS goals. Developing reliable and interpretable metrics specifically designed for evaluating application tasks from a 3D Gaussian perspective is essential to track progress and guide future research. Existing metrics mainly evaluate accuracy after projection onto 2D images, lacking meaningful insights into the inherent 3D geometric and semantic consistency of generated or edited scenes.

### v3

@@ 4.1 Performance Benchmarking: 3DGS Segmentation
TABLE II: Quantitative 3D interactive segmentation on NVOS [276] and SPIn-NeRF[67] in terms of mIoU and mAcc.
@@
TABLE III: Performance comparison of open-vocabulary 3D semantic segmentation on ScanNet [244] benchmark. The evaluation metrics are mIoU and mAcc, respectively.
@@
$\bullet$ Results. Following the taxonomy in Sec.2.1.1, we report results across four representative segmentation settings. For 3D instance segmentation (Table I), Unified-Lift [20] achieves the best overall performance on both the Replica [275] and LERF-Mask [22] datasets, demonstrating the effectiveness of consistency-aware end-to-end learning. In the setting of 3D interactive segmentation (Table II), GaussianCut [29] performs best on NVOS [243], while Click-Gaussian [24] attains the highest mIoU on SPIn-NeRF [67]. Notably, GaussianCut is a training-free method that flexibly supports both 2D clicks and text prompts as input. For open-vocabulary 3D semantic segmentation (Table III), PanoGS [118] and InstanceGaussian [31] achieve the best performance under point-based evaluation protocols, whereas CAGS [108] performs best under 3D Gaussian-based evaluation. In the open-vocabulary 2D semantic segmentation setting, VLGaussian [28] reaches a state-of-the-art mIoU of 97.1% on the 3D-OVS benchmark [63], significantly outperforming existing approaches in Table IV. Table V further presents results for two subtasks: open-vocabulary 2D localization (detection) and semantic segmentation. FMLGS [102] achieves the best performance on the segmentation task, while VLGaussian [28] leads in localization. Across these settings, most methods leverage powerful vision-language foundation models such as CLIP and SAM to enhance open-vocabulary capabilities.
TABLE IV: Quantitative open-vocabulary 2D semantic segmentation on 3D-OVS [63] dataset in terms of mIoU.
@@
TABLE V: Quantitative results on LERF-OVS [26] Dataset. We report the open-vocabulary localization accuracy (mAcc) and 2D semantic segmentation (mIoU).
@@
LangSplatV2[104] [NeurIPS’25] 51.8 74.7 56.4 82.1 72.2 93.2 59.1 86.4
TABLE VI: Performance comparison of 3DGS editing on Mip-NeRF360 [245] and IN2N [75] datasets. The similarity for CLIP Text-Image Direction,CLIP Text-Image, CLIP Image-Image, DINO are denoted as $\text{CLIP}_{dir}$, $\text{CLIP}_{T2I}$, $\text{CLIP}_{I2I}$, and $\text{DINO}_{sim}$, respectively.
@@
TABLE VII: Performance comparison of 3DGS style transfer on Mip-NeRF360 [245] dataset, evaluated by LPIPS and RMSE.
@@ 4.2 Performance Benchmarking: 3DGS Editing
$\bullet$ Datasets. We summarize the key features of the datasets in Appendix Table A.2, with detailed descriptions provided below.
@@
$\bullet$ Results. For 3D image editing tasks, the most commonly used datasets are Mip-NeRF360 [245] and IN2N [75]. We adopt these two datasets as benchmarks to record the performance of various methods. However, as most existing 3D editing methods are evaluated on different scene subsets, it remains difficult to perform a comprehensive and fair comparison under a unified protocol. Given that IN2N [75] is widely adopted as a baseline, we report the relative gains of each method over IN2N to enable an indirect yet relatively fair comparison, as summarized in Table VI. Overall, GaussianVTON[170] demonstrates the best performance across multiple metrics. For the style transfer task, we use Mip-NeRF 360 [245] as the benchmark and compare the results of different methods in Table VII, where SemanticSplatSty.[182] achieves the best overall performance.
TABLE VIII: Quantitative 3DGS generation on GSO [247] dataset. CLIP Image-Image Similarity as $\text{CLIP}_{I2I}$.
@@
TABLE IX: Quantitative 3DGS generation on Objaverse [246] dataset. CLIP Text-Image Similarity as $\text{CLIP}_{T2I}$.
@@ 4.3 Performance Benchmarking: 3DGS Generation
$\bullet$ Datasets. We summarize the key features of the datasets in Appendix Table A.3, with detailed descriptions provided below.
@@
Objaverse [246] is a large-scale dataset of objects with 800K+ 3D models with descriptive captions, tags, and animations.
@@
$\bullet$ Results. We select two representative datasets, GSO [247] and Objaverse [246], as benchmarks for 3DGS generation. The performance of various methods is summarized in Table VIII and Table IX, respectively. Due to the limited number of publicly comparable baselines, we provide a quantitative assessment based on the available results. On GSO dataset, NovelGS [235] achieves the best performance by tailoring the diffusion process to suit the 3D Gaussian representation. It supports both text-to-3D and image-to-3D generation, demonstrating strong generalization across input modalities. On Objaverse dataset, Atlas-Gaussians [225] performs best, leveraging latent space optimization to enhance generation quality. It primarily supports text-conditioned generation and shows promising results on open-domain content.
@@ 5 Future Directions
$\bullet$ Faithful Evaluation Metrics. Current evaluation protocols often inherit metrics from traditional 2D or NeRF-based tasks, which may not align well with 3DGS goals. Developing reliable and interpretable metrics specifically designed for evaluating application tasks from a 3D Gaussian perspective is essential to track progress and guide future research. Existing metrics mainly evaluate accuracy after projection onto 2D images, lacking meaningful insights into the inherent 3D geometric and semantic consistency.
@@ 6 Conclusion
Appendices
A.1 Dataset Statistics
Due to space limitations in the main manuscript, we present the full table of the statistics of representative 3DGS segmentation, editing, and generation datasets in Table A.1, Table A.2, and Table A.3.
TABLE A.1: Statistics of representative 3DGS segmentation datasets. See §4.1 for more detailed descriptions.
  Dataset Venue #Scene #Views(Avg.) Characterization
TABLE A.2: Statistics of representative 3DGS editing datasets. See §4.2 for more detailed descriptions.
  Dataset Venue #Scenes #Views(Avg.) Characterization
TABLE A.3: Statistics of representative 3DGS generation datasets. See §4.3 for more detailed descriptions.
  Dataset Venue #Type #Scenes Characterization
[…]
A.2 Functional Application Tasks
A.2.1 Human Avatar
A.2.2 SLAM